CYCLOPOLYMERIZATION OF ACETYLENE TO BENZYNE AND NAPHTHALENE

\underline{DILRUKSHI HEWAGE}, RUCHIRA SILVA AND DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Reactions of acetylene (C_{2}H_{2}) with laser-vaporized La atoms produced La(C_{6}H_{4}) and La (C_{10}H_{8}) in supersonic molecular beams. The organic fragments in these complexes were benzyne and naphthalene. The benzyne species was produced by the La-mediated cyclotrimerization of three acetylene molecules, whereas naphthalene was formed likely by the cyclization of the transient benzyne with two additional acetylene molecules. These cyclized products were identified by mass-analyzed threshold ionization mass spectroscopy, which measured adiabatic ionization energies and several vibrational frequencies. The measured ionization energies were 40875 (5) cm^{−1} for La(C_{6}H_{4}) and 36767 (5) cm^{−1} for La(C_{10}H_{8}). The most active vibrational transitions of both complexes were metal-ligand stretching with 326 cm^{−1} for La(C_{6}H_{4}) and 286 cm^{−1} for La (C_{10}H_{8}). By combining the spectra with theoretical calculations, the ground electronic states of the neutral complexes were determined to be ^{2}A_{1} (C_{2v}) and ^{2}A′ (C_{s}) and those of the corresponding ions were ^{1}A_{1} (C_{2v}) and ^{1}A′ (C_{s}) for La (C_{6}H_{4}) and La(C_{10}H_{8}), respectively.